# Compute-Constrained Data Selection

### v2

@@ D.1 Finetuning Settings
We follow standard practices in LLM finetuning Wang et al. (2023a); Ivison et al. (2023) and use the AdamW optimizer with beta-parameters $(\beta_{1},\beta_{2})=(0.9,0.99)$.
@@ D.3 Training Datasets
For training, we use the same four processed datasets as in Wang et al. (2023a); Ivison et al. (2023), all of which are either human-authored or annotated. Details are provided in Table 8. The FLAN V2 and COT datasets are derived from existing NLP datasets, while DOLLY and OPEN ASSISTANT 1 contain open-ended generation examples with human-written responses. These datasets vary in format, sequence length, and tasks. Following Wang et al. (2023a), we standardize their format using the ‘Tulu’ structure.
@@
Table 8: Details of training dataset from Wang et al. (2023a). Len. is short for token length.
@@ Appendix F Additional Results: IFEval
Appendix G Additional Analysis: Multiple Task-Specific Models.
@@
Figure 10: Multiple Task-Specific Model Break-Even Analysis for 7B BBH. Performance under compute-constraints reach the finetuned Pareto frontier at 30 tasks.
@@
Figure 12: Multiple Task-Specific Model Break-Even Analysis for 7B IFEval. Performance under compute-constraints reach the finetuned Pareto frontier at 2 tasks, surpassing it at 5, 10 tasks.
@@
Appendix H Additional Analysis: Data Similarity.

### v4

@@ D.1 Finetuning Settings
We follow standard practices in LLM finetuning Wang et al. (2023); Ivison et al. (2023) and use the AdamW optimizer with beta-parameters $(\beta_{1},\beta_{2})=(0.9,0.99)$.
@@ D.3 Training Datasets
For training, we use the same four processed datasets as in Wang et al. (2023); Ivison et al. (2023), all of which are either human-authored or annotated. Details are provided in Table 8. The FLAN V2 and COT datasets are derived from existing NLP datasets, while DOLLY and OPEN ASSISTANT 1 contain open-ended generation examples with human-written responses. These datasets vary in format, sequence length, and tasks. Following Wang et al. (2023), we standardize their format using the ‘Tulu’ structure.
D.4 Evaluation Datasets
We evaluate our method on three benchmark datasets: MMLU (Hendrycks et al., 2020), BBH (Suzgun et al., 2022), and IFEval (Zhou et al., 2023). Table 9 contains more details about each tasks. Each subtask comes with few-shot examples or sample responses, which are used as validation set $\mathcal{V}$ for data selection and as few-shot in-context learning demonstration in evaluation.
An important aspect of the data selection approach is the size of the validation set used for computing utility scores. While a small validation set might cause overfitting or insufficient representation of the target task, our experiments indicate that even modestly sized validation sets (e.g., containing more than 50 examples) are sufficient for methods like BM25 and Embed to perform effectively. Since we are selecting large subsets from the training data, the precision required from the validation set is not overly stringent. These methods can effectively capture relevant training samples without significant risk of overfitting, suggesting that, in practice, reasonably sized validation sets suffice for similarity-based data selection in LLM finetuning.
@@ D.4 Evaluation Datasets
Table 8: Details of training dataset from Wang et al. (2023). Len. is short for token length.
Dataset # Shot # Tasks $|\bm{\mathcal{V}}|$ $|\bm{\mathcal{T}}|$ Answer Type
MMLU 5 57 285 18,721 Letter options
BBH 3 23 69 920 COT and answer
IFEval 1 - 50 541 Open Generation
Table 9: Statistics of evaluation datasets. The selection of evaluation tasks cover different kinds of answer types.
@@ Appendix F Additional Results: IFEval
Appendix G Extrapolating from the Parametric Function
In Figure 9 and Figure 10, we extrapolate the parametric fits obtained from our 7B and 13B model runs by increasing the ratio between the training model size and the data selection model size, thereby effectively reducing the proportional cost of gradient-based data selection. Graphically, this is seen when the parametric fit crosses the existing compute-optimal frontier—a tipping point at which the method becomes compute-optimal.
We found that the training-to-selection model size ratio should be approximately 5 for perplexity-based data selection, and 10 for gradient-based data selection. This suggests that using more expensive data selection methods becomes advantageous when the training model size largely exceeds the selection model size by the compute-optimal ratio. Under these conditions, meaningful efficiency gains can be achieved compared to cheaper data selection methods.
Figure 9: Extrapolating Compute-Optimal Model Size for MMLU. By extrapolating from the parametric fits obtained from the 7B and 13B model results for MMLU, we can find the compute-optimal ratio between the training model size and the selector model size required for the perplexity-based and gradient-based method. For perplexity-based data selection, our extrapolation suggests that training model should be larger (5x) than the selector model—around 35B model parameters. For gradient-based data selection to be compute-optimal, our extrapolation suggest that the training model should be significantly larger (10x) than the selector model—specifically, around 70B parameters in this case.
Figure 10: Extrapolating Compute-Optimal Model Size for BBH. We perform a similar extrapolation for BBH and find results consistent with those in Figure 9. Specifically, for the perplexity-based method to be compute-optimal, the training model should be 5x larger than the selection model, and for the gradient-based method, the training model should be approximately 10x larger than the selection model.
Appendix H Additional Analysis: Multiple Task-Specific Models.
@@ Appendix G Extrapolating from the Parametric Function
Figure 12: Multiple Task-Specific Model Break-Even Analysis for 7B BBH. Performance under compute-constraints reach the finetuned Pareto frontier at 10 tasks.
@@
Figure 14: Multiple Task-Specific Model Break-Even Analysis for 7B IFEval. Performance under compute-constraints reach the finetuned Pareto frontier at 5 tasks, surpassing it at 10 tasks.
@@
Appendix I Additional Analysis: Data Similarity